# Generation of High Order Harmonics in Vacuum for Various Configurations of Interacting Electromagnetic Field

### v1

@@ V.2 Generation of the high-order harmonics by two intersecting Gaussian beams
$\times\left\{{\color[rgb]{1,0,0}2}n\biggl{[}\zeta\tilde{\Theta}_{\zeta}+i\upsilon\tilde{\Theta}_{\upsilon}-i\left(1-\frac{\upsilon^{2}+\zeta^{2}}{2}\right)+\tilde{\Theta}_{\upsilon}^{2}\biggr{]}+3\left[2i-(1+i)(\upsilon^{2}+\zeta^{2})+\frac{3}{4}\upsilon^{2}\right]\right\}$
@@
$\tilde{\Theta}_{\upsilon}=W\Theta\cos\phi$ and $\tilde{\Theta}_{\zeta}=W\Theta\sin\phi$. Details of the derivations of Eqs. (94) and (95), are presented in Secs. C.4 and C.5 of App. C. Calculation of the integral in Eq. (95) comes down to several calculations of the integrals of the type: $\int_{-\infty}^{\infty}u^{\ell}\exp(au-bu^{2})\,du$ for integer $\ell$. As a result, we may obtain explicit analytic expressions for $\tilde{\mathcal{K}}_{n}$ for any integer $n$. Such expressions for $n=1$, 2, 3 and 4 are presented in App. C.5.
@@ VI Summary and Discussion
A general theory of the HOH generation obtained in the frame of the Heisenberg-Euler Electrodynamics is presented in this paper. We obtain general expressions for the matrix elements (47)-(49) of the nonlinear current, entering the expression (40) for angular distribution of the HOH intensity for monochromatic $in$-coming wave. The case of the ‘same’ polarization is only considered in this paper. This limitation simplifies considerably the final expressions, and takes into account an observation that typically mixed polarization leads to diminishing of the HOH generation rate [40]. We apply these general expressions for the two cases: i) for the $4\pi$ dipole $in$-coming field that gives maximal efficiency for the HOH generation at given total power of the $in$-coming wave; and ii) for the case of two crossing Gaussian beams. These results are summarized in i) Eq. (59) and Tab. 2, and in ii) Eq. (97) and Tab. 3, respectively. Angular distributions of the HOH radiation are presented in Figs. 4 and 7, respectively.
Power of the HOH of the order $m$, generated by the mechanism considered in this paper, is proportional to $P^{m+2}$, where $P$ is total power of the $in$-coming wave. However, we cannot increase the power unlimitedly. At a sufficiently high power, the wave will produce electron-positron plasma [64, 9, 37, 52, 53, 54, 55, 56, 57] as an avalanche triggered by the Schwinger mechanism [38, 39, 50, 51] of the electron-positron pair production. Radiation of such a plasma will lead to screening of the observable effect of HOH generation in vacuum. Thus, total rate of the HOH generation in vacuum that can be investigated experimentally is bounded from above. We considered this effect and the results are summarized in Tab. 2 separately for the $E1$ and $M1$ mode for the $4\pi$-dipole $in$-coming wave. We see that it would be hardly to detect the harmonics with the number $m$ higher than 5 for optical lasers.
Our general expressions (40) and (47)-(49) are obtained in lowest order of perturbation theory relative to $in$-coming wave power. They corresponds to the process
@@
if we describe it in terms of elementary plane waves. Here the wave 4-vectors, $k_{i=1,2,\dots,2n+3}$, belong the $in$-coming wave (beams), whereas the 4-wave vector, $\mathcal{K}$, corresponds to the generated HOH beam. All $k_{i}$ have the same frequency $\omega$, whereas the time component of the vector $\mathcal{K}$ corresponds to the frequency $\Omega=(2n+1)\omega$ of the HOH.
@@
However, the matrix element of this process vanishes exactly due to the structure of the initial Heisenberg-Euler Lagrangian, described by Eqs. (1) and (4). Taking into account vanishing photon mass in vacuum, the conservation law corresponding to the reaction (109) can be fulfilled only when $\boldsymbol{k}_{i}\parallel\boldsymbol{\mathcal{K}}$ that leads to vanishing of the the product of the invariant bi-linear forms $f$ and $g$ (9)-(10), entering Eq. (42). This exact cancellation was not taken into account in Ref. [46], when the relevant integrals evaluated approximately. It leads to considerable overestimation of effectiveness of the 3rd order harmonic generation. As a result, the lowest order process giving contribution to the process of HOH generation is described by Eqs. (40) and (47)-(35), corresponding to the process (108), as considered in Secs. III.2.
@@
will not contain at all the small parameter $W^{-2}$ in its leading order. A particular case of the process (110) was considered in Ref. [40]. We showed in Sec. V.3 that if the 2nd beam power is less than a critical value, defined in Eq. (107), then the process (108) is more probable than the process (110), regardless of the focusing degree, $W^{-1}$, while it is small. However, the processes (110), considered actually in Ref. [40], dominates over (108), when the power of the 2nd incident laser beam becomes higher than the critical value under the same other limitations. This conclusion concerns only not very high order, $m=2n+1$, of the harmonics: $n\ll W^{2}$. Really, this condition will be not so restrictive.
@@ B.1 General expressions
$\times\biggl{\{}-j_{1}(\omega r)\left[\sin\theta\cos\varphi{\color[rgb]{1,0,0}-}\frac{2i}{\omega r}\left(\cos\Theta\cos^{2}\theta{\color[rgb]{1,0,0}+}\sin\Theta\sin\theta\cos\theta\cos\varphi\right)\right]$
$\frac{i}{\omega r}\left(\partial_{r}(rj_{1}(\omega r))\right)\left[\cos\Theta\sin^{2}\theta{\color[rgb]{1,0,0}-}\sin\Theta\cos\theta\sin\theta\cos\varphi\right]\biggl{\}}$ (118)
@@
$H_{n}={\color[rgb]{1,0,0}-}\int f^{n}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}$
${\color[rgb]{1,0,0}-}2\pi\int\limits_{0}^{\infty}r^{2}\,dr\int\limits_{0}^{\pi}\sin\theta\,d\theta\,f^{n}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,\overline{f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)}\,$, (122)
$K_{n}={\color[rgb]{1,0,0}-}\int f^{n+1}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}={\color[rgb]{1,0,0}-}2\pi\int\limits_{0}^{\infty}r^{2}\,dr\int\limits_{0}^{\pi}\sin\theta\,d\theta\,f^{n+1}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,\overline{f\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)}\,.$ (123)
@@
Combining the expressions  (122) and (123) and expressions (116), (117), (120) and (121) for the invariant $f$, and using the definitions (57) and (58), and recovering the amplitude multiplier $E_{0}$ in Eqs. (51)-(55), we obtain the following explicit expressions for $\bar{H}_{n}(\Theta)$ and $\bar{K}_{n}(\Theta)$:
$\bar{H}_{n}(\Theta)={\color[rgb]{1,0,0}-}2\int\limits_{0}^{\infty}\rho^{2}\,d\rho\int\limits_{0}^{\pi/2}\sin\theta\,d\theta\,\left[j_{1}^{2}(\rho)\sin^{2}\theta+\frac{1}{\rho^{2}}\left[\partial_{\rho}\left(\rho j_{1}(\rho)\right)\right]^{2}\sin^{2}\theta+\frac{4}{\rho^{2}}j_{1}^{2}(\rho)\cos^{2}\theta\right]^{n}$
@@
$\bar{K}_{n}(\Theta)={\color[rgb]{1,0,0}-}2\int\limits_{0}^{\infty}\rho^{2}\,d\rho\int\limits_{0}^{\pi/2}\sin\theta\,d\theta\,\left[j_{1}^{2}(\rho)\sin^{2}\theta+\frac{1}{\rho^{2}}\left[\partial_{\rho}\left(\rho j_{1}(\rho)\right)\right]^{2}\sin^{2}\theta+\frac{4}{\rho^{2}}j_{1}^{2}(\rho)\cos^{2}\theta\right]^{n+1}$

### v2

@@ V.2 Generation of the high-order harmonics by two intersecting Gaussian beams
$\tilde{\Theta}_{\upsilon}=W\Theta\cos\phi$ and $\tilde{\Theta}_{\zeta}=W\Theta\sin\phi$. Details of the derivations of Eqs. (98) and (99), are presented in Secs. C.4 and C.5 of App. C. Calculation of the integral in Eq. (99) comes down to several calculations of the integrals of the type: $\int_{-\infty}^{\infty}u^{\ell}\exp(au-bu^{2})\,du$ for integer $\ell$. As a result, we may obtain explicit analytic expressions for $\tilde{\mathcal{K}}_{n}$ for any integer $n$. Such expressions for $n=1$, 2, 3 and 4 are presented in App. C.5. These explicit analytic expressions prove the fact that the matrix elements $H_{n}$ and $K_{n}$, introduced in Eqs. (49) and (50) do not generally vanish for this incident beams geometry. This is an example for the general statement considered in Sec. III.2.
@@
The HOH generation process considered in this section, whose rate is described by Eq. (101), can be described in terms of Eq. (52) in the following way. The $(2n+1)$ wave 3-vectors $\boldsymbol{k}_{1,2,\dots,2n+1}$ belong to the beam No 1, propagating mainly in the $x$-direction, whereas the 3-vectors $\boldsymbol{k}_{2n+2,\,2n+3}$ belong to the beam No 2, propagating mainly in $z$-direction. See Fig. 5 for the better visualization. The 3-vectors from the first group have typically small angles, $\sim 1/W$, between each other. The same is valid for the wave vectors from the second group. As a result, the matrix elements $H_{n}$ and $K_{n}$ have an additional smallness proportional to a power of $1/W$, as it was considered above. This gives the factor $1/W^{4n}$ that has appeared finally in Eq. (101).
@@
Eq. (101) is the exact expression in the following sense. We separate the lowest (leading) order of the small parameters $E/E_{S}$ in an implied general expression for the intensity of the harmonic, then separate the leading order with respect to the small parameter $\alpha$, and then we separate the leading order with respect to $1/W$.
The next order terms in the perturbation theory with respect to $\mathcal{L}_{HE}$ in Eq. (1) or with respect to the term $\alpha Q$ in Eq. (14) will contain higher powers of $\alpha$ and, probably, lower powers of $1/W$. In this case, we should demand a relationship between $\alpha$ and $1/W$ for Eq. (102) would remain the valid leading order for the harmonic generation rate. Our preliminary estimations for the cases $n=1$, 2 and 3 show that the condition looks like: $\alpha W^{2}\ll 1$. We assume that this condition can be considered as a necessary condition for any $n\geq 1$. We see that Eq. (101) contains rather high powers of the small parameter $1/W$. For this reason this condition does not look as so restrictive from the practical point of view. A counterplay between the small parameters $E/E_{S}$ and $1/W$ is considered below in Sec. V.3.
@@ VI Summary and Discussion
A general theory of the HOH generation obtained in the frame of the Heisenberg-Euler Electrodynamics in vacuum is presented in this paper. The HE electrodynamics is based on existence of two small parameters: $\alpha$, the fine structure constant, and $E/E_{S}$, where $E$ is a typical absolute value of the electric field strength in the focus of the incident EM wave, and $E_{S}$ is the QED critical field. We obtain the general expression (51) for the angular distribution of the HOH intensity for the case of monochromatic $in$-coming wave. The case of the ‘same’ polarization is only considered in this paper. This limitation simplifies considerably the final expressions, and takes into account an observation that typically mixed polarization leads to diminishing of the HOH generation rate [41]. Our general expression (51) is an exact leading order term of a general expression in the following sense. We keep only the lowest powers of the small parameters $E/E_{S}$ and $\alpha$.
We apply these general expressions for the two cases: i) for the $4\pi$ dipole $in$-coming field that gives maximal efficiency for the HOH generation at given total power of the $in$-coming wave; and ii) for the case of two crossing Gaussian beams. These results are summarized in i) Eq. (63) and Tab. 2, and in ii) Eq. (101) and Tab. 3, respectively. Angular distributions of the HOH radiation are presented in Figs. 4 and 7, respectively. The expressions for the case (ii) contain an additional small parameter, $1/W$, that characterizes a small convergence angle of the Gaussian beams. A counterplay of this small parameter with $E/E_{S}$ and $\alpha$ is considered in Secs. V.2 and V.3.
Power of the HOH of the order $m$, generated by the mechanism considered in this paper, is proportional to $P^{m+2}$, where $P$ is total power of the $in$-coming wave. However, we cannot increase the power unlimitedly. At a sufficiently high power, the wave will produce electron-positron plasma [67, 10, 38, 54, 55, 56, 57, 58, 59, 60] as an avalanche triggered by the Schwinger mechanism [39, 40, 52, 53] of the electron-positron pair production. Radiation of such a plasma will lead to screening of the observable effect of HOH generation in vacuum. Thus, total rate of the HOH generation in vacuum that can be investigated experimentally is bounded from above. We considered this effect and the results are summarized in Tab. 2 separately for the $E1$ and $M1$ mode for the $4\pi$-dipole $in$-coming wave. We see that it would be hardly to detect the harmonics with the number $m$ higher than 5 for optical lasers.
Our general expression (51), containing the matrix elements (49) and (50), is obtained in the lowest order of our perturbation theory relative to $in$-coming wave power. They corresponds to the process
@@
if we describe it in terms of elementary plane waves. Here the wave 4-vectors, $k_{i=1,2,\dots,2n+3}$, belong to the $in$-coming wave (beams), whereas the 4-wave vector, $\mathcal{K}$, corresponds to the generated HOH beam. All $k_{i}$ have the same frequency $\omega$, whereas the time component of the vector $\mathcal{K}$ corresponds to the frequency $\Omega=(2n+1)\omega$ of the HOH.
@@
However, the matrix element of this process vanishes exactly due to the structure of the initial Heisenberg-Euler Lagrangian, described by Eqs. (1) and (4). Taking into account vanishing photon mass in vacuum, the conservation law corresponding to the reaction (113) can be fulfilled only when $\boldsymbol{k}_{i}\parallel\boldsymbol{\mathcal{K}}$ that leads to vanishing of the the product of the invariant bi-linear forms $f$ and $g$ (9)-(10), entering Eq. (43). This exact cancellation was not taken into account in Ref. [47], when the relevant integrals evaluated approximately. It leads to considerable overestimation of effectiveness of the 3rd order harmonic generation. As a result, the lowest order process giving contribution to the process of HOH generation is described by Eqs. (41) and (48)-(36), corresponding to the process (112), as considered in Sec. III.2.
@@
will not contain at all the small parameter $W^{-2}$ in its leading order. A particular case of the process (114) was considered in Ref. [41]. We showed in Sec. V.3 that if the 2nd beam power is less than a critical value, defined in Eq. (111), then the process (112) is more probable than the process (114), regardless of the focusing degree, $W^{-1}$, while it is small. However, the processes (114), considered actually in Ref. [41], dominates over (112), when the power of the 2nd incident laser beam becomes higher than the critical value under the same other limitations.
@@ B.1 General expressions
$\times\biggl\{-j_{1}(\omega r)\left[\sin\theta\cos\varphi-\frac{2i}{\omega r}\left(\cos\Theta\cos^{2}\theta+\sin\Theta\sin\theta\cos\theta\cos\varphi\right)\right]$
$\frac{i}{\omega r}\left(\partial_{r}(rj_{1}(\omega r))\right)\left[\cos\Theta\sin^{2}\theta-\sin\Theta\cos\theta\sin\theta\cos\varphi\right]\biggl\}$ (122)
@@
$H_{n}=-\int f^{n}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}$
$2\pi\int\limits_{0}^{\infty}r^{2}\,dr\int\limits_{0}^{\pi}\sin\theta\,d\theta\,f^{n}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in\,*}\right)\,\overline{f\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)}\,$, (126)
$K_{n}=-\int f^{n+1}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,f\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)\,d\boldsymbol{x}=-2\pi\int\limits_{0}^{\infty}r^{2}\,dr\int\limits_{0}^{\pi}\sin\theta\,d\theta\,f^{n+1}\left(\tilde{\boldsymbol{A}}^{in},\tilde{\boldsymbol{A}}^{in}\right)\,\overline{f\left(\tilde{\boldsymbol{A}}^{in\,*},\tilde{\boldsymbol{\mathcal{A}}}^{out\,*}\right)}\,.$ (127)
@@
Combining the expressions  (126) and (127) and expressions (120), (121), (124) and (125) for the invariant $f$, using the definitions (61) and (62), and recovering the amplitude multiplier $E_{0}$ in Eqs. (55)-(59), we obtain the following explicit expressions for $\bar{H}_{n}(\Theta)$ and $\bar{K}_{n}(\Theta)$:
$\bar{H}_{n}(\Theta)=-2\int\limits_{0}^{\infty}\rho^{2}\,d\rho\int\limits_{0}^{\pi/2}\sin\theta\,d\theta\,\left[j_{1}^{2}(\rho)\sin^{2}\theta+\frac{1}{\rho^{2}}\left[\partial_{\rho}\left(\rho j_{1}(\rho)\right)\right]^{2}\sin^{2}\theta+\frac{4}{\rho^{2}}j_{1}^{2}(\rho)\cos^{2}\theta\right]^{n}$
@@
$\bar{K}_{n}(\Theta)=-2\int\limits_{0}^{\infty}\rho^{2}\,d\rho\int\limits_{0}^{\pi/2}\sin\theta\,d\theta\,\left[j_{1}^{2}(\rho)\sin^{2}\theta+\frac{1}{\rho^{2}}\left[\partial_{\rho}\left(\rho j_{1}(\rho)\right)\right]^{2}\sin^{2}\theta+\frac{4}{\rho^{2}}j_{1}^{2}(\rho)\cos^{2}\theta\right]^{n+1}$